Week of 180521

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday: Whit Monday holiday

  • The meeting will be held on Tuesday instead.

Tuesday

Attendance:

  • local: Borja (Monit), Charles (Databases), Julia (WLCG), Maarten (ALICE, WLCG), Vincent (Security)
  • remote: Andrzej (ATLAS), Christoph (CMS), Christian (NDGF), Darren (RAL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Overall smooth running
    • Production jobs
      • Production running fine at 315k jobs and good contributions from HPC
      • Slightly increased rates of jobs failing, but sites reacting to tickets even during weekend holidays
      • Recent special productions: completion of data17 reprocessing and production of derivations mostly completed

Andrzej mentioned for the completion of data17 it was needed to get extending data from tapes and thanked the tape team.

  • CMS reports ( raw view) -
    • Average CPU usage: ~170k cores for production and ~50k cores for analysis
    • Site Status Board (SSB)not reachable early last week, GGUS:135095
      • SSB is rather critical for CMS computing operations
      • Some authentication issue for CMS supporter seems to remain: GGUS:135179
    • Had a small campaign to bring sites to most recent FTS client 3.7.8
      • Older clients seem to have trouble with most recent servers
      • Requested to have this also added to CVMFS repo grid.cern.ch, GGUS:135134
    • CNAF CMS storage developed a problem Monday (May 21st ) evening: GGUS:135249 and GGUS:135145
      • Meanwhile fixed by restarting GridFTP doors
    • CMS VOMS server became unresponsive: GGUS:135257
    • Some trouble with CMS tests and Singularity
      • Required changes are applied
      • False alarms at sites are being corrected from availability statistics

Christoph highlighted the not so fast user support for the SSB related tickets. CVMFS request to be finished soon.

  • ALICE -
    • NTR

Sites / Services round table:

  • ASGC: nc
  • BNL: nc
  • CNAF: nc
  • EGI: nc
  • FNAL: nc
  • IN2P3: NTR
  • JINR: nc
  • KISTI: nc
  • KIT: nc
  • NDGF: NTR
  • NL-T1: nc
  • NRC-KI: nc
  • OSG: nc
  • PIC: nc
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services: nc
  • CERN storage services: nc
  • CERN databases:
    • Performance issues affecting LHCB due to the update of their queries, making them to take more time. (reverted to previous version while investigation is ongoing).
  • GGUS: NTR
  • Monitoring:
    • Final reports for the April 2018 availability sent around
  • MW Officer: nc
  • Networks: NTR
  • Security: Security incident at a non-EGI academic site:
    • Please check for outbound network traffic to the IPs mentioned in the broadcast
    • One of the IP seems to be a scanner, expect to see (benign) inbound scans

AOB:

Edit | Attach | Watch | Print version | History: r11 < r10 < r9 < r8 < r7 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r11 - 2018-05-22 - BorjaGarridoBear
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback